OpenStudy (anonymous):

One solution of the equation 6x^2-19x+3=0 is 1/6 What is the other solution?

OpenStudy (anonymous):

You can try doing the quadratic formula.

OpenStudy (lgbasallote):

if 1/6 is a solution that means x = 1/6 is a root that means 6x - 1 is a factor so just divide 6x^2 - 19x + 3 by 6x - 1

OpenStudy (anonymous):

How do I do that?

OpenStudy (lgbasallote):

long division? synthetic division? isnt that your lesosn?

OpenStudy (anonymous):

sum of the roots of Quadratic equation ax^2+bx+c=0 is -b/a so in our case a=6 b=-19 so the sum of roots should be ,let unknown root is r 1/6+r=-(-19)/6 \[\Large r+\frac{1}{6}=\frac{19}{6}\] OR \[\Large r=\frac{19}{6}-\frac{1}{6}\] simplify it and you got your answer !!

OpenStudy (lgbasallote):

lol

OpenStudy (anonymous):

yeah !!!!

OpenStudy (lgbasallote):

wonder why wolfram disagrees though

OpenStudy (anonymous):

let me check my Algebra then .

OpenStudy (lgbasallote):

lol jk

OpenStudy (anonymous):

It is \[Sum = -\frac ba = \frac{19}{6}\]

OpenStudy (anonymous):

So, you suntract 1/6 to get 18/6 which simplifies...

OpenStudy (anonymous):

lol i was right !!!!!!!

OpenStudy (anonymous):

Yup XD Just always remember the sign rule for these types of things ;)

OpenStudy (anonymous):

@lgbasallote wolfram agrees with me :P

OpenStudy (lgbasallote):

didnt you see the lol jk? lol

OpenStudy (anonymous):

oops!!! now i saw lol jk :P

OpenStudy (anonymous):

Sorry guys my computer froze!!

OpenStudy (anonymous):

So the other solution is 18/6 ?!?

OpenStudy (anonymous):

yes 18/6=3

OpenStudy (anonymous):

Well spankyou. Spankyou all very much.